# APCTP-Pre2023-010

### v2

@@ III.2 Broad turns
The last term is differentiated by $x$ and is hence first-order in terms of $x^{-1}$. By using this transformation we can reduce the problem to a zeroth-order diagonal part plus a perturbation term, and our goal is to consistently solve the zeroth-order part only. This, however, does not mean we can discard entirely this perturbation term. Since the matrix $U$ is not uniquely defined, ignoring the first-order part in the previous equation and finding $Y$ to zeroth order will make $X$ ambiguous as it depends on the specific choice of $U$. To completely remove this ambiguity, we note that for $U$ we have the freedom to multiply each of its columns by an arbitrary function, which amounts to rescaling the eigenvectors of $J_{(0)}$. Thus we can choose $U$ in such a way that the subdominant term in (82) has only non-diagonal elements. This choice can be shown to be
$\tilde{U}=U\exp\left(\int\,\mathrm{d}x\,P\right)\,,\,$ (83)
where $P$ denotes the diagonal elements of the first order part $U^{-1}J_{(1)}U-U^{-1}U_{,x}$ which yields the solution of our original problem as  [^6] In a sense, to derive the solution in the so called “adiabatic picture” we apply the interaction picture not on the original system $X^{\prime}=J\,X$ but on the “rotated system” described by $Y=U^{-1}\,X$. Then we break the new stability matrix in two parts, the first being $J_{\rm D}+P$ which we can solve analytically (since it consists of diagonal terms), and the second being the interaction part consisting of the matrix with entries the off-diagonal elements of $U^{-1}J_{(1)}UY-U^{-1}U_{,x}$ (see also PhysRevA.45.3329 ; Blanes:2008xlr ).
$X(N)=U\exp\left[\int\,\mathrm{d}N\left(J_{\rm D}+P\right)\right]\,.\,$ (84)
The integral of the second term is related to the Berry’ phase Berry:1984jv .
For our previous example we find the eigenvectors of $J_{(0)}$ as $e_{1}=(-i,1)^{T}$, $e_{2}=(i,1)^{T}$ and the diagonal higher-order terms of elements of (82) as $\text{diag}(1/x,1/x)$. Therefore, by redefining the matrix $U$ as
$\tilde{U}=\begin{pmatrix}-ix&ix\\ x&x\end{pmatrix}\,,$ (85)
@@
$F_{Y}=\text{diag}\left(e^{ix},e^{-ix}\right)\,,$ (86)
and subsequently the fundamental matrix of our original problem via $F_{X}=\tilde{U}F_{Y}$. Therefore, we find the general solution for the curvature perturbation and its momentum as
$\mathcal{R}$ $=c_{1}xe^{ix}+c_{2}xe^{-ix}\,,$ (87)
@@
which are accurate up to $\mathcal{O}(1)$ corrections. It should also be clear that the Hubble friction terms provide the correct linear dependence on $x$, even though they are suppressed in the $x^{-1}$ expansion. This happens because they contribute to the $\log x$ corrections on the final result and thus cannot be consistently ignored in the final result even to leading order (see e.g. Gong:2001he for a derivation in a similar spirit). The result obtained with the previous method coincides with the WKB solution when considering an asymptotic expansion $\mathcal{R}=\exp\left(x\mathcal{R}_{(0)}+\mathcal{R}_{(1)}+\cdots\right)$ and keeping only the first two terms. Although for the single-field case the standard WKB method is simpler, for more fields substituting the above asymptotic expressions would result into a coupled time-dependent system considerably more involved than what we obtain with the method described so far.
Moving to the two-field case, the $4\times 4$ system in the $x$-parameterization becomes
$\mathcal{R}_{,x}$ $=p_{\mathcal{R}}-{2\Omega\over x}\mathcal{S}\,,$ (88)
$p_{\mathcal{R},x}$ $={2\over x}p_{\mathcal{R}}-\mathcal{R}\,,$ (89)
$\mathcal{S}_{,x}$ $=p_{\mathcal{S}}\,,$ (90)
$p_{\mathcal{S},x}$ $={2\Omega\over x}p_{\mathcal{R}}+{2\over x}p_{\mathcal{S}}-\left(1+{m_{\rm eff}\over x^{2}}\right)\mathcal{S}\,.$ (91)
This, however, is not in the appropriate form because the turn rate introduces a new expansion parameter. Assuming that the turn rate is large ($\Omega\gg 1$) we will soon see that there are effectively three regimes of interest:
  1. 1.
    $x\gg\Omega$, where both the Hubble friction and the coupling terms are subdominant;
  2. 2.
    $x\sim\Omega$, which is roughly the period when most of the amplification takes place;
  3. 3.
    $x\ll\Omega$, which is the regime of validity of the effective single-field description.
@@
$\mathcal{R}_{,\theta}$ $=\tilde{p}_{\mathcal{R}}+\mathcal{S}\,,$ (92)
$\tilde{p}_{\mathcal{R},\theta}$ $=-{3\over 2\Omega}\tilde{p}_{\mathcal{R}}+\left({x\over 2\Omega}\right)^{2}\mathcal{R}\,,$ (93)
$\mathcal{S}_{,\theta}$ $=p_{\mathcal{S}}\,,$ (94)
$p_{\mathcal{S},\theta}$ $=-\tilde{p}_{\mathcal{R}}-{3\over 2\Omega}p_{\mathcal{S}}-\left[\left({x\over 2\Omega}\right)^{2}+{m_{\rm eff}\over 4\Omega^{2}}\right]\mathcal{S}\,,$ (95)
where $\tilde{p}_{\mathcal{R}}\equiv P_{\mathcal{R}}/(2\Omega)$. We are interested in solving the system for $x<2\Omega$ and we observe that the terms quadratic in $x$ become order $1/\Omega$ when $x\sim\sqrt{\Omega}$. Having terms of the wrong order in the definition of $J_{(0)}$ may affect the domain of validity of derived approximate solutions; this point will soon become clear. In those cases when separating the different terms we actually have to further divide this regime in two more: $x\sim(\Omega,\sqrt{\Omega})$ and $x\sim(\sqrt{\Omega},1)$. For these two regimes the zeroth- and first-order stability matrices become
$J_{(0)}\equiv\begin{pmatrix}0&1&1&0\\ -\left(\dfrac{x}{2\Omega}\right)^{2}&0&0&0\\ 0&0&0&1\\ 0&-1&-\left[\left(\dfrac{x}{2\Omega}\right)^{2}+\dfrac{m_{\rm eff}}{4\Omega^{2}}\right]&0\end{pmatrix}\quad\text{and}\quad J_{(1)}\equiv\begin{pmatrix}0&0&0&0\\ 0&-\dfrac{3}{2\Omega}&0&0\\ 0&0&0&0\\ 0&0&0&-\dfrac{3}{2\Omega}\end{pmatrix}\,,$ (96)
for the former case and
$\tilde{J}_{(0)}\equiv\begin{pmatrix}0&1&1&0\\ 0&0&0&0\\ 0&0&0&1\\ 0&-1&-\dfrac{m_{\rm eff}}{4\Omega^{2}}&0\end{pmatrix}\quad\text{and}\quad\tilde{J}_{(1)}\equiv\begin{pmatrix}0&0&0&0\\ -\left(\dfrac{x}{2\Omega}\right)^{2}&-\dfrac{3}{2\Omega}&0&0\\ 0&0&0&0\\ 0&0&-\left(\dfrac{x}{2\Omega}\right)^{2}&-\dfrac{3}{2\Omega}\end{pmatrix}\,,$ (97)
for the latter case. First, we focus on the scales $x\sim\Omega$: The four eigenvalues of $J_{(0)}$ are {±$\{\pm\lambda_{\rm im},\pm\lambda_{\rm p}\}$ with
$\lambda_{\rm im}$ $\equiv{i\over 2\Omega\sqrt{2}}\sqrt{2x^{2}+m_{\rm eff}+\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}\,,$ (98)
$\lambda_{\rm p}$ $\equiv{1\over 2\Omega\sqrt{2}}\sqrt{-2x^{2}-m_{\rm eff}+\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}\,,$ (99)
where the subscript “im” reflects the fact that the first two eigenvalues are always imaginary, and $\lambda_{\rm p}$ becomes positive for $x^{2}<4\Omega^{2}-m_{\rm eff}$. Following the same steps as before, we find the general solution for $\mathcal{R}$ and $\mathcal{S}$ as
$\mathcal{R}$ $=xf_{+}(x)\left(c_{1}e^{-\int\lambda_{\rm im}}+c_{2}e^{\int\lambda_{\rm im}}\right)+xf_{-}(x)\left(c_{3}e^{-\int\lambda_{\rm p}}+c_{4}e^{\int\lambda_{\rm p}}\right)\,,$ (100)
$\mathcal{S}$ $=xf_{+}(x)s_{\rm im}\left(c_{1}e^{-\int\lambda_{\rm im}}-c_{2}e^{\int\lambda_{\rm im}}\right)+xf_{-}(x)s_{\rm p}(x)\left(-c_{3}e^{-\int\lambda_{\rm p}}+c_{4}e^{\int\lambda_{\rm p}}\right)\,,$ (101)
@@
$f_{\pm}(x)$ $\equiv\left[1-{4\Omega^{2}\pm m_{\rm eff}\over\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}-{4\Omega^{2}(2x^{2}\mp m_{\rm eff})\over m_{\rm eff}^{2}+16x^{2}\Omega^{2}}\right]^{1/4}\,,$ (102)
$s_{{\rm im}({\rm p})}(x)$ $\equiv-{8\Omega^{2}\lambda_{{\rm im}({\rm p})}^{2}+2x^{2}\over 8\Omega^{2}\lambda_{{\rm im}({\rm p})}}\,.$ (103)
The solution is ill-behaved for $x\to\sqrt{4\Omega^{2}-m_{\rm eff}}$ which is the point where the matrix $U$ becomes non-diagonalizable (the analogue of a turning point in the WKB approximation). This happens because the eigenvalues ±$\pm\lambda_{p}$ switch from imaginary to positive/negative. To properly account for this one has to calculate the Jordan form of $J$ to find the approximate solution (the analogue of the Airy functions). However, we can circumvent this by assuming that the two perturbations are still in the Bunch Davies vacuum shortly after $x$ becomes smaller than $\sqrt{4\Omega^{2}-m_{\rm eff}}$ and match the perturbations and their derivatives at that smaller scale with the previous two analytic expressions.
Except for the massless case $m_{\rm eff}=0$ for which the solution remains accurate until horizon crossing, for all other masses with $m_{\rm eff}>0$, the solution loses accuracy as $x$ departs from the value $x_{*}=\sqrt{2\Omega}$. This happens because certain off-diagonal terms of (82) become order $\Omega^{-1/2}$ on these scales and hence our approximation becomes worse, and which as we explained earlier is due to not separating consistently terms of different order. Only for the massless case $m_{\rm eff}=0$ the off-diagonal terms remain of order $\mathcal{O}(\Omega^{-1})$ up to horizon crossing and hence (100) and (101) remain accurate (see the left panel of figure 2), whereas for the masses $m_{\rm eff}\leq 4\Omega^{2}$ one has to also consider the first correction in the formula (84). For the borderline case $m_{\rm eff}=4\Omega^{2}$, which divides the models for which $\lambda_{\rm p}$ can become positive from those where $\lambda_{\rm p}$ remains a complex number, the off-diagonal terms become order one at the scale $\sqrt{\Omega}$ and, hence, after this scale the system needs to be solved using the matrices (97). By doing so and expanding in powers of $\Omega^{-1}$ we find that at $x\sim\sqrt{\Omega}$ the amplitudes of the two perturbations are almost equal $|\mathcal{R}|\sim|\mathcal{S}|\sim\Omega^{1/4}$, and after that the curvature perturbation freezes $|\mathcal{R}|\sim|\tilde{p}_{\mathcal{R}}|_{x=\sqrt{\Omega}}\sim(2\Omega)^{1/4}$ (recovering the result of Cremonini:2010ua ), whereas the isocurvature perturbation decays as $x^{3/2}/\sqrt{2\Omega}$ (see the right panel of figure 2).
Figure 2: Left: The ratio of isocurvature to curvature perturbations for the massless case with $\Omega=10$. Soon after $\lambda_{\rm p}$ becomes positive this ratio is dominated by $s_{\rm p}$. Right: The curvature and isocurvature power spectra for the borderline case $m_{\rm eff}=4\Omega^{2}$ in the regime $x<\sqrt{2\Omega}$ for $\Omega=20$ (solid), 50 (dashed) and 100 (dotted). In both cases $N=0$ corresponds to horizon crossing.
@@
Away from this point these functions are almost constant so we can regularize them e.g. by replacing $f_{-}=(\cdots)^{1/4}\rightarrow(i+\cdots)^{1/4}$. Moreover, the constants $c_{i}$ can be found by matching (100),(101) with the values of the two perturbations before the point $x_{0}$, e.g. the Bunch-Davies solutions. By doing so we find that at horizon crossing $|c_{i}|$ yield $\mathcal{O}(1)$ corrections so we can ignore them and quote the final result for the curvature perturbation:
@@
$\mathcal{P}_{\mathcal{R}}\approx\mathcal{P}_{0}|f_{-}|^{2}\left.\exp\left(2\int_{\sqrt{4\Omega^{2}-m_{\rm eff}}}^{1}\,\mathrm{d}x\lambda_{\rm p}\right)\right|_{x=1}\,,$ (104)
@@
with $\mathcal{P}_{0}\approx x^{2}$ for $x\approx 1$. Note that extending the upper integration limit in the above integral to $x\rightarrow 0$ yields the amplification factor $\pi\Omega\left(2-\sqrt{m_{\rm eff}}/\Omega\right)$ that was mentioned in Bjorkmo:2019qno .
IV Criteria for the growth of perturbations

### v3

@@ 3.2 Broad turns
The last term is differentiated by $x$ and is hence first-order in terms of $x^{-1}$. By using this transformation we can reduce the problem to a zeroth-order diagonal part plus a perturbation term, and our goal is to consistently solve the zeroth-order part only. This, however, does not mean we can discard entirely this perturbation term. Since the matrix $U$ is not uniquely defined, ignoring the first-order part in the previous equation and finding $Y$ to zeroth order will make $X$ ambiguous as it depends on the specific choice of $U$. To completely remove this ambiguity, we note that for $U$ we have the freedom to multiply each of its columns by an arbitrary function, which amounts to rescaling the eigenvectors of $J_{(0)}$. Thus we can choose $U$ in such a way that the subdominant term in (3.30) has only non-diagonal elements. This choice can be shown to be
$\tilde{U}=U\exp\left(\int\,\mathrm{d}x\,P_{\rm D}\right)\,,\,$ (3.31)
where $P_{\rm D}$ denotes the diagonal elements of the first-order part in (3.30), i.e. we split it as follows:
$U^{-1}J_{(1)}UY-U^{-1}U_{,x}=P_{\rm D}+P_{\rm od}\,.\,$ (3.32)
Based on that observation we rewrite (3.30) as a sum of a diagonal and an off-diagonal part:
$Y_{,x}=D_{\rm Y}Y+P_{\rm od}Y\,,\,$ (3.33)
with $D_{\rm Y}\equiv J_{\rm D}+P_{\rm D}$, which yields the leading-order solution for the fundamental matrix of our original problem as [^6] In a sense, to derive the solution in the so called “adiabatic picture” we apply the interaction picture not on the original system $X^{\prime}=J\,X$ but on the “rotated system” described by $Y=U^{-1}\,X$. Then we break the new stability matrix in two parts, the first being $J_{\rm D}+P_{\rm D}$ which we can solve analytically (since it consists of diagonal terms), and the second being the interaction part consisting of the matrix with entries the off-diagonal elements of $U^{-1}J_{(1)}UY-U^{-1}U_{,x}$ (see also [63, 64]).
$F_{\rm X}=U\exp\left[\int\,\mathrm{d}ND_{\rm Y}\right]\,.\,$ (3.34)
Notice that $\int\,\mathrm{d}NP_{\rm D}$ is related to the Berry’s phase [65].
For our previous example we find the eigenvectors of $J_{(0)}$ as $e_{1}=(-i,1)^{T}$, $e_{2}=(i,1)^{T}$ and the diagonal higher-order terms of elements of (3.30) as $\text{diag}(1/x,1/x)$. Therefore, by redefining the matrix $U$ as
$\tilde{U}=\begin{pmatrix}-ix&ix\\ x&x\end{pmatrix}\,,$ (3.35)
@@
$F_{\rm Y}=\text{diag}\left(e^{ix},e^{-ix}\right)\,,$ (3.36)
and subsequently the fundamental matrix of our original problem via $F_{\rm X}=\tilde{U}F_{\rm Y}$. Therefore, we find the general solution for the curvature perturbation and its momentum as
$\mathcal{R}$ $=c_{1}xe^{ix}+c_{2}xe^{-ix}\,,$ (3.37)
@@
which are accurate up to $\mathcal{O}(1)$ corrections. It should also be clear that the Hubble friction terms provide the correct linear dependence on $x$, even though they are suppressed in the $x^{-1}$ expansion. This happens because they contribute $\log x$ corrections and thus cannot be consistently ignored in the final result even to leading order (see e.g. [57] for a derivation in a similar spirit). The result obtained with the previous method coincides with the WKB solution when considering an asymptotic expansion $\mathcal{R}=\exp\left(x\mathcal{R}_{(0)}+\mathcal{R}_{(1)}+\cdots\right)$ and keeping only the first two terms. Although for the single-field case the standard WKB method is simpler, for more fields substituting the above asymptotic expressions would result into a coupled time-dependent system considerably more involved than what we obtain with the method described so far.
Moving to the two-field case, we observe that the $4\times 4$ system (3.1)-(3.4) includes three scales, the Hubble scale $x$, the turn rate $\Omega$ and the mass parameter $m_{\rm eff}$. Again, we are interested in the subhorizon regime, as on superhorizon scales one can apply the transfer functions formalism to study the evolution of perturbations (see e.g. [15, 66, 67, 68, 69]). Therefore, we first write the system in the $x$-parameterization:
$\mathcal{R}_{,x}$ $=p_{\mathcal{R}}-{2\Omega\over x}\mathcal{S}\,,$ (3.38)
$p_{\mathcal{R},x}$ $={2\over x}p_{\mathcal{R}}-\mathcal{R}\,,$ (3.39)
$\mathcal{S}_{,x}$ $=p_{\mathcal{S}}\,,$ (3.40)
$p_{\mathcal{S},x}$ $={2\Omega\over x}p_{\mathcal{R}}+{2\over x}p_{\mathcal{S}}-\left(1+{m_{\rm eff}\over x^{2}}\right)\mathcal{S}\,.$ (3.41)
The previous set of equations describe a perturbative expansion in inverse powers of $x$ which remains valid until $x$ becomes of order $\text{max}(\sqrt{m_{\rm eff}},2\Omega,1)$, where we also assume $\Omega\gg 1$. Note that during the first phase the two perturbations decay as almost massless fields. Depending on the relative magnitude of $2\Omega$ to $\sqrt{m_{\rm eff}}$ there are two subsequent types of behaviour. The first case includes models with $m_{\rm eff}\geq 4\Omega^{2}$ which have been studied in early works of effective field theory descriptions, e.g. [56, 61]. For these models the amplification of the power spectrum as a function of the turn rate scales at most as $\mathcal{P}_{\mathcal{R}}\sim\sqrt{\Omega}$ and hence they are not appropriate candidates for PBH formation (see also Appendix B). For this reason for the remainder of this section we focus on the latter case, namely $m_{\rm eff}<4\Omega^{2}$.
@@
$\mathcal{R}_{,\theta}$ $=\tilde{p}_{\mathcal{R}}+\mathcal{S}\,,$ (3.42)
$\tilde{p}_{\mathcal{R},\theta}$ $=-{3\over 2\Omega}\tilde{p}_{\mathcal{R}}-\left({x\over 2\Omega}\right)^{2}\mathcal{R}\,,$ (3.43)
$\mathcal{S}_{,\theta}$ $=p_{\mathcal{S}}\,,$ (3.44)
$p_{\mathcal{S},\theta}$ $=-\tilde{p}_{\mathcal{R}}-{3\over 2\Omega}p_{\mathcal{S}}-\left[\left({x\over 2\Omega}\right)^{2}+{m_{\rm eff}\over 4\Omega^{2}}\right]\mathcal{S}\,,$ (3.45)
where $\tilde{p}_{\mathcal{R}}\equiv P_{\mathcal{R}}/(2\Omega)$. Similarly to the single-field case we split the full matrix to zeroth- and first-order parts:
$J_{(0)}\equiv\begin{pmatrix}0&1&1&0\\ -\left(\dfrac{x}{2\Omega}\right)^{2}&0&0&0\\ 0&0&0&1\\ 0&-1&-\left[\left(\dfrac{x}{2\Omega}\right)^{2}+\dfrac{m_{\rm eff}}{4\Omega^{2}}\right]&0\end{pmatrix}\quad\text{and}\quad J_{(1)}\equiv\begin{pmatrix}0&0&0&0\\ 0&-\dfrac{3}{2\Omega}&0&0\\ 0&0&0&0\\ 0&0&0&-\dfrac{3}{2\Omega}\end{pmatrix}\,.$ (3.46)
The four eigenvalues of $J_{(0)}$ are {±$\{\pm\lambda_{\rm im},\pm\lambda_{\rm p}\}$ with
$\lambda_{\rm im}$ $\equiv{i\over 2\Omega\sqrt{2}}\sqrt{2x^{2}+m_{\rm eff}+\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}\,,$ (3.47)
$\lambda_{\rm p}$ $\equiv{1\over 2\Omega\sqrt{2}}\sqrt{-2x^{2}-m_{\rm eff}+\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}\,,$ (3.48)
where the subscript “im” reflects the fact that this eigenvalue is always imaginary, whereas $\lambda_{\rm p}$ becomes positive for $x^{2}<4\Omega^{2}-m_{\rm eff}$. Following the same steps as before, we find the general solution for $\mathcal{R}$ and $\mathcal{S}$ as
$\mathcal{R}$ $=xf_{+}(x)\left(c_{1}e^{-\int\lambda_{\rm im}}+c_{2}e^{\int\lambda_{\rm im}}\right)+xf_{-}(x)\left(c_{3}e^{-\int\lambda_{\rm p}}+c_{4}e^{\int\lambda_{\rm p}}\right)\,,$ (3.49)
$\mathcal{S}$ $=xf_{+}(x)s_{\rm im}\left(c_{1}e^{-\int\lambda_{\rm im}}-c_{2}e^{\int\lambda_{\rm im}}\right)+xf_{-}(x)s_{\rm p}(x)\left(-c_{3}e^{-\int\lambda_{\rm p}}+c_{4}e^{\int\lambda_{\rm p}}\right)\,,$ (3.50)
@@
$f_{\pm}(x)$ $\equiv\left[1-{4\Omega^{2}\pm m_{\rm eff}\over\sqrt{m_{\rm eff}^{2}+16x^{2}\Omega^{2}}}-{4\Omega^{2}(2x^{2}\mp m_{\rm eff})\over m_{\rm eff}^{2}+16x^{2}\Omega^{2}}\right]^{1/4}\,,$ (3.51)
$s_{{\rm im}({\rm p})}(x)$ $\equiv-{8\Omega^{2}\lambda_{{\rm im}({\rm p})}^{2}+2x^{2}\over 8\Omega^{2}\lambda_{{\rm im}({\rm p})}}\,.$ (3.52)
The previous solution has been derived under the assumption of negligible off-diagonal terms in (3.30) for the rotated vector $Y$. Thus, we need to check whether these terms remain small compared to one.
First, we focus on the massless case $m_{\rm eff}=0$. The off-diagonal elements $P_{\rm od}$ remain $\mathcal{O}(\Omega^{-1})$ until horizon crossing, hence we can trust the expansion up to that order. However, $P_{\rm od}$ diverge as $x\rightarrow 2\Omega$ which is the point where the eigenvalues ±$\pm\lambda_{\rm p}$ switch from imaginary to positive/negative, indicating that $J_{(0)}$ becomes non-diagonalizable at that point (the analogue of a turning point in the WKB approximation). This further implies that there will be a small region around $x=2\Omega$ where the approximate solution has to be found with a different method. Nevertheless, it suffices to consider the two solutions as Bunch-Davies modes and match the two expressions at a later time. Moving to masses satisfying $0<m_{\rm eff}<4\Omega^{2}$ the range of validity of the above approximate solution is more limited. The off-diagonal elements diverge as $x\rightarrow\sqrt{4\Omega^{2}-m_{\rm eff}}$, which again is the point where $\lambda_{\rm p}$ crosses zero. In addition, as $x\rightarrow\sqrt{2\Omega}$ the off-diagonal terms become $\mathcal{O}(\Omega^{-1/2})$ and eventually scale as $\sqrt{4\Omega^{2}}/\sqrt{4\Omega^{2}-m_{\rm eff}}$ when $x\rightarrow 1$. This means that the approximate solution (3.50) is strictly valid for $x\in\left(\sqrt{2\Omega},x_{*}\right)$ with $x_{*}<\sqrt{4\Omega^{2}-m_{\rm eff}}$. For the scales $x\in(1,\sqrt{2\Omega})$ a more accurate version would require a different split of zeroth and first-order terms or another rotation $X=UY$, where now $U$ diagonalizes a different matrix. However, simply using the previous formulae we find a deviation of at most a factor of 10 closer to the borderline limit. Specifically for $m_{\rm eff}=4\Omega^{2}$ the analytical solution completely fails (Appendix B shows how to obtain the correct solution in this case).
Figure 2: Numerical (solid) versus analytic approximation (dashed) for the curvature (blue solid-orange dashed) and isocurvature (black solid-red dashed) power spectra (normalized by the single-field expression $\mathcal{P}_{0}$). Even though we have considered moderate values for the turn rate the approximation remains accurate up to horizon crossing ($N=0$). We have checked that larger values of the turn rate improve the accuracy of the analytic approximation. We consider in the left panel the massless case $m_{\rm eff}=0$ with $\Omega=8$, and in the right panel $m_{\rm eff}=2\Omega^{2}$ with $\Omega=11$.
@@
Regarding the constants $c_{i}$ these can be found by matching (3.49) and (3.50) with the values of the two perturbations before the point $x_{*}$, e.g. the Bunch-Davies solutions. For our purposes it is sufficient to treat them as $\mathcal{O}(1)$ numbers and hence at horizon crossing we quote the final result for the curvature perturbation:
@@
$\left.\mathcal{P}_{\mathcal{R}}\right|_{x=1}\approx\mathcal{P}_{0}|f_{-}|^{2}\left.\exp\left[2\int_{\sqrt{4\Omega^{2}-m_{\rm eff}}}^{1}\,\mathrm{d}x\left(-{2\Omega\over x}\right)\lambda_{\rm p}\right]\right|_{x=1}\,,$ (3.53)
@@
with $\mathcal{P}_{0}\approx x^{2}$ for $x\approx 1$. Note that extending the upper integration limit in the above integral to $x\rightarrow 0$ yields the exponential amplification factor $\pi\Omega\left(2-\sqrt{m_{\rm eff}}/\Omega\right)$ that was mentioned in [59]. We have plotted the approximate solution in figure 2 for two choices of the mass parameter and by considering $c_{3}=c_{4}=1$ and $c_{1}=c_{2}=0$.
4 Criteria for the growth of perturbations